OpenStudy (celinegirl):

The table below shows two equations: Equation 1: |2x − 3| + 5 = 4 Equation 2: |5x + 3| − 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution, and equation 2 has solutions x = 2, −3.2. The solutions to equation 1 are x = 1, 2, and equation 2 has no solution. The solutions to equation 1 are x = 1, 2, and equation 2 has solutions x = 2, −3.2.

OpenStudy (solomonzelman):

General example; \(\color{#000000 }{ \displaystyle |ax-b|=c }\) \(\color{#000000 }{ \displaystyle ax-b=c }\) or \(\color{#000000 }{ \displaystyle ax-b=-c }\) \(\color{#000000 }{ \displaystyle ax=b+c }\) or \(\color{#000000 }{ \displaystyle ax=b-c }\) So the answer would be: \(\color{#000000 }{ \displaystyle x=(b+c)/a }\) or \(\color{#000000 }{ \displaystyle x=(b-c)/a }\)

OpenStudy (celinegirl):

confused... :/

OpenStudy (solomonzelman):

Do you know what absolute value is?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle |x| }\) - "absolute value of x".